Ask
your own question, for FREE!
Mathematics
6 Online
can someone check my work??? please!!! You are watching your favorite soccer player on TV. Statistics show that he has a 37% chance of scoring a goal each time he shoots. He shoots twice. What is the probability that he scores a goal both times? steps please!!!
Still Need Help?
Join the QuestionCove community and study together with friends!
37/100 * 37/100 = 1369/10000 1369/10000 1369 divided by 1000 o.1369
Do you know if I used the correct steps to solve this?
@amistre64 can you help me with the steps please
@terenzreignz can you help me?
Quite correct ^_^ The two events are independent, since scoring a goal the first time will not affect whether or not the player scores the next (ideally :P ) So you simply multiply the probabilities.
Still Need Help?
Join the QuestionCove community and study together with friends!
THANKS!!!!
:)
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
7 hours ago
1 Reply
0 Medals
4 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
2 days ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal